OpenStudy (anonymous):

(x^2-3x+2)-(3x^2-5x-1) Answer please thanks

OpenStudy (anonymous):

hey gurl ;D

OpenStudy (katrinakaif):

-2x^2 +2x +1 Hey! =)

OpenStudy (anonymous):

awe, i got it wrong ? I got -3x^4+14x^3-20x^2+7x+2 am i wrong??

OpenStudy (chaise):

Would it not be +3? +2 - -1?

OpenStudy (katrinakaif):

Your right Chaise! Sorry again =( UTK fresh all your supposed to do is substract..dont multiply

OpenStudy (chaise):

It's fine, we all make mistakes.

OpenStudy (anonymous):

Oh fck lol thanks so much ;D

OpenStudy (katrinakaif):

No problem =)

OpenStudy (anonymous):

im sorry i still don't get it step by step please thanks sorry

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

(x^2)-3x+2-(3x^2)+5x+1=0 (-2x^2)+2x+3=0 (-2x^2)+2x=3 -2((x^2)-x)=3 (x^2)-x=-3/2 (x^2)-x+3/2=0 Right here I would use the quadratic formula x= (-b+-Sqrt((b^2)-4ac)/2a a=1, b=-1, c=3/2 x=1+-sqrt((1^2)-4(1)(3/2))/2(1) x=(1+-sqrt(6))/2 That is pretty much your answer without using imaginary numbers

OpenStudy (phi):

Is this multiplying or subtraction?? if it is multiplying (stuff) times -(more stuff) you did it right. if it is subtraction, (stuff) minus (more stuff) see Kat's answer See http://www.wolframalpha.com/input/?i=%28x%5E2-3x%2B2%29*-%283x%5E2-5x-1%29

OpenStudy (katrinakaif):

(x^2-3x+2)-(3x^2-5x-1) So first we have to distribute the NEGATIVE to (3x^2 - 5x -1) When we do that we get: -3x^2 + 5x +1 Now we collect like terms: (x^2-3x+2 - 3x^2 + 5x + 1) Collect the x^2s first, by addicng their coeffiencts. 1x^2 - 3x^2 = 6x^2 Now add the coefficents of the x terms only: -3x + 5x = 2x Now collect the like terms of constants: 2 + 1 = 3 PUt it all together and get : -2x^2 + 2x + 3